OpenStudy (anonymous):

Jessica has some markers. She has 5 more black markers than blue markers. She has 4 times as many red markers as black markers. She has 8 blue markers. How many markers does Jessica have altogether?

OpenStudy (solomonzelman):

blue + 5 = black red = black × 4 if blue =8, then blue + 5 = black 8+ 5 = black 13 = black And, red = 13 × 4 = 52 red+blue+black = 52 + 8 + 13 = ? you tell me.... :)

OpenStudy (anonymous):

73 thank you

OpenStudy (solomonzelman):

You welcome !